# Equal Contribution   Equal Advising https://sirui-xu.github.io/InterDreamer/

### v1

@@ C Additional Details of Experimental Setup
Figure A: We use Amazon Mechanical Turk [96] to build an annotation platform. We provide instructions to guide the annotator to split a long sequence into several short sub-sequences with their start and end frames, and then annotate each sub-sequence. We inform annotators that our collected data are used for text-motion generation when they accept this work.
Datasets. We have included a screenshot of our annotation platform in Fig. A. Upon acceptance of our work, we plan to update the Acknowledgements section with the identities of the annotators. Our annotations are further diversified by the GPT-4 [67]. The prompt used for this purpose is: I’m going to give you  an essay, and I want you to give me three sentences in English of
varying degrees of complexity, like the following example: ‘‘A
person lifts something with both hands, dumps it to the right, and
then puts it back down.’’ A person reaches to the left for
something, then reaches to the right as if to dump it, and then
puts it back to the left. A person picks up something with both
hands, tilts it, and then puts it down. To make the text more
complex, you can add more detailed adjectives and adverbs and
complicate sentence structure and verbs. The input text is ‘‘A man
leaned his upper right thigh against a small table and stretched
his left hand and left foot outward.’’ Please give me three texts
that vary in complexity but keep the meaning of the sentence the
same.
@@
Metrics. In Sec. 4.1, we introduce the metrics employed in this paper. This section details the formula for the metric CMD proposed. The formulations for other metrics are available in the existing literature [116, 25]. CMD quantifies the discrepancy between the contact maps of ground truth interactions and those synthesized one. In this context, a contact map is characterized by the proportion of time each body part $\{\boldsymbol{p}_{i}\}_{i=1}^{P}$ maintains active contact. Here, $\boldsymbol{p}_{i}$ denotes the percentage of time during which the body part $i$ is less than a threshold distance from the object. And the metric is defined as,
$CMD=\frac{1}{P}\sum_{i=1}^{P}\|\boldsymbol{p}_{i}-\boldsymbol{p}_{i}^{\mathrm{GT}}\|_{1},$ (14)
@@
where $\boldsymbol{p}_{i}^{\mathrm{GT}}$ is from the ground truth contact map, $P$ is the number of the body parts defined in the SMPL [58], and we define the distance threshold as $0.03$ m.
@@
Implementation Details. The segment in the MDP contains $m=4$ frames. The dynamics model, which includes 2 dynamics blocks as described in the main paper, is trained on the BEHAVE training set [7], with a batch size of 32, a latent dimension of 64, and for 500 epochs. For rollout after the initial step $t>1$, our dynamics model is trained to predict over a longer timeframe ($F=3\times m=12$), exceeding the past motion duration ($H=m=4$). For the initial step $t=1$, we train a separate dynamics model to forecast a duration of $F=15$ given the past motion over $H=1$ frame, consistent with Sec. B.2. The optimization process is conducted over $300$ epochs, utilizing a learning rate of $0.01$.
@@ D Additional Qualtitative Results
User Study. We conduct a double-blind user study. Given two annotations, we use our dynamics model with InterDiff [116] and with our proposed vertex control to generate 8 samples respectively. We design pairwise evaluations. Considering one pair from ours and another one from InterDiff, human judges are asked to determine which interaction has the better interaction quality. From the results of 11 human evaluations in Table A, our approach has a success rate of 61.4% against the baseline.
Figure B: Qualitative results from the interaction retrieval. We demonstrate that our proposed handcraft interaction retrieval (unified color person) and learning-based interaction retrieval (textured person, where textures are from [12]) can extract diverse and realistic interactions.
Interaction Retrieval. In addition to the results in the main paper and the demo video, we here visualize the intermediate retrieval results. Fig. B depicts both the handcraft retrieval and learning-based retrieval procedure, resulting in a diverse set of interactions that are both high-quality and semantically aligned.
Table A: User study on the zero-shot text-to-interaction task. Under pairwise human voting results, our generated results significantly outperform the baseline considering the interaction fidelity.
Method w/ InterDiff [116] vs. w/ vertex control (ours)
w/ InterDiff [116] vs. N/A 38.6%
w/ vertex control (ours) 61.4% N/A
E Potential Negative Societal Impact
Some potential negative societal impacts include: (i) Our approach can be used to synthesize realistic human motion interacting with objects, which could potentially lead to the creation of misinformation. (ii) Our approach evaluates on real behavioral information, which may raise privacy concerns. However, our model utilizes a processed representation (SMPL [58]) of the human motion that retains minimal identifying details, in contrast to raw data or images. This aspect can be positively regarded as a feature that enhances privacy.

### v2

@@ C Additional Details of Experimental Setup
Figure A: We use Amazon Mechanical Turk [120] to build an annotation platform. We provide instructions to guide the annotator to split a long sequence into several short sub-sequences with their start and end frames, and then annotate each sub-sequence. We inform annotators that our collected data are used for text-motion generation when they accept the job.
Figure B: Full log of our high-level planning.
Figure C: Qualitative results from the interaction retrieval. We demonstrate that our learning-based interaction retrieval can extract diverse and realistic interactions.
Datasets. We include a screenshot of our annotation platform in Figure A. Our annotations are further diversified by GPT-4 [87]. The prompt used for this purpose is: I’m going to give you a description, and I would like to have three rewritten sentences with varying degrees of complexity, following the example: ‘‘...’’ The input text is ‘‘...’’ Please give me three texts that vary in complexity but keep the meaning of the sentence the same. This results in (i) less complexity: someone holds a backpack and steps left; (ii) middle complexity: a person holds a backpack in front of them with both hands and takes a step to the left; (iii) more complexity: with both hands, a person clutches a heavy backpack firmly and brings it close to their body, then steps to the left with their left leg.
@@
Metrics. In Sec. 4.1, we introduce the metrics employed in this paper. This section details the formula for the proposed metric CMD. The formulations for other metrics are available in the existing literature [148, 34]. CMD quantifies the discrepancy between the contact maps of ground truth interactions and those synthesized one. In this context, a contact map is characterized by the proportion of time $\{\bm{p}_{i}\}_{i=1}^{P}$ each body part maintains active contact. Here, $\bm{p}_{i}$ denotes the percentage of time during which the body part $i$ is less than a threshold distance from the object. And the metric is defined as,
$\mathrm{CMD}=\frac{1}{P}\sum_{i=1}^{P}\|\bm{p}_{i}-\bm{p}_{i}^{\mathrm{GT}}\|_{1},$ (11)
@@
where $\bm{p}_{i}^{\mathrm{GT}}$ is from the ground truth contact map, $P$ is the number of body parts defined in SMPL [76], and we set the distance threshold as $0.03$ m.
@@
Implementation Details. The segment in the MDP contains $m=4$ frames. The dynamics model, which includes 2 dynamics blocks as described in the main paper, is trained on the BEHAVE training set [7], with a batch size of 32, a latent dimension of 64, and for 500 epochs. For rollout after the initial step $t>1$, our dynamics model is trained to predict over a longer timeframe ($F=3\times m=12$), exceeding the past motion duration ($H=m=4$). For the initial step $t=1$, we train a separate dynamics model to forecast a duration of $F=15$ given the past motion over $H=1$ frame, consistent with Sec. B.2. The optimization process is conducted over $300$ epochs, utilizing a learning rate of $0.01$. The dynamic model is trained on an NVIDIA A40 GPU for a day. Our full log for high-level planning is presented in Figure B.
@@ D Additional Qualtitative Results
Interaction Retrieval. We here visualize the intermediate retrieval results. Figure C depicts the results from learning-based retrieval, resulting in a diverse set of interactions that are both high-quality and semantically aligned.
Qualitative Experiments on OMOMO dataset. Figure D exemplifies our method that is able to generalize effectively to the OMOMO [66] dataset, despite our dynamics model not being trained on its object geometry or annotations.
Figure D: Qualitative results on the OMOMO [66] dataset. Our method generalizes well on the OMOMO objects and annotations unseen in training. Frames are separately visualized. Here, our synergized models are GPT-4 [87] and MotionGPT [46].
E Failure Cases
We present failure cases of our method in our website, consisting of (i) inconsistency of interaction with textual description, (ii) inconsistency of human actions with textual descriptions, and (iii) wrong object category inferred by LLM.
F Potential Negative Societal Impact
Some potential negative societal impacts include: (i) Our approach can synthesize realistic human motion interacting with objects, which could be misused to create deceptive or harmful content, such as portraying individuals in false situations. This could contribute to the spread of misinformation. (ii) Our method evaluates real behavioral information, raising potential privacy concerns. Although our model utilizes a processed representation (SMPL [76]) of human motion that retains minimal identifying details – unlike raw data or images – its ability to simulate human-object interactions could still be exploited for unauthorized surveillance or behavioral analysis. For instance, with photorealistic textures, it might be used to model and generate personal habits or movements without consent, posing risks of privacy violations. However, the use of a processed representation can be positively viewed as a privacy-enhancing feature, as it minimizes the exposure of personally identifiable details.